OpenStudy (anonymous):

an african cheetah runs at a speed of 120 km/h what is the speed in feet per second \

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 @morgen

jaynator495 (jaynator495):

GEt outa here im taking this onje :P

jaynator495 (jaynator495):

\(~~~~~~~~\sf\Huge\rlap{\color{Cyan}{Welcome}}{\color{blue}{\;Welcome}}\rlap{\color{teal}{To}}{\color{#0044ff}{\;To}}\rlap{\color{#4700FF}{OpenStudy!}}{\color{#A300FF}{\;OpenStudy!}}\\\sf\Huge\rlap{\color{#FF0000}{We}}{\color{#FF7300}{\;We}}\rlap{\color{#FFCF00}{Have}}{\color{#FFFF00}{\;Have}}\rlap{\color{#D1FF00}{Been}}{\color{#75FF00}{\;Been}}\rlap{\color{#00FFA8}{Expecting}}{\color{#00E6FF}{\;Expecting}}\rlap{\color{#00A1FF}{You!}}{\color{#0045FF}{\;You!}}\) \(\bullet\) Scince You Are New To Openstudy, please Click On This Link And Read The CoC http://openstudy.com/code-of-conduct/ This Will Tell You The Rules Of Openstudy, If You Dont Follow The Rules You Will Get In Trouble \(\bullet\) Scince You Are New To Openstudy, please Click On This Link And Read The TaC http://openstudy.com/terms-and-conditions This Is Legal Stuff... Its Not AS Important But Its Still Important None The Less \(\bullet\) Please Click This Link To Learn LaTeX https://openstudy.com/study#/updates/54ee31cae4b0dc552b64ba51 With LaTeX you can make Amazing Text Art... You Can Use Hyper Links... If You Think Of It... Most Likely You Can Do It, With This Tutorial Unleash You Inner Creativity And Learn Latex... PS Don't Forget To Medal! \(\bullet\) If You Need Anything Feel Free To Tag Me And I'll Do My Best Help Out

TheSmartOne (thesmartone):

We need to use conversions to get your answer, so: \(\sf\Large \frac{120 km}{1 hr}\times \frac{1hr}{60 mins}\times \frac{1 min}{60 secs}\) And that gives you a unit of km per hour. because the hours and the minutes in the numerator and denominator cancel out leaving us with seconds.

TheSmartOne (thesmartone):

And then going further to make it into feet, \(\sf\Large \frac{120 km}{1 hr}\times \frac{1hr}{60 mins}\times \frac{1 min}{60 secs}\times\frac{1 km}{3280.84ft}=?\) Just multiply the numbers all the numbers and at the end put the units as \(\sf\Large\frac{ft}{secs}\)

TheSmartOne (thesmartone):

@morgen